what is 2a^2-30+108=0 rounded to the nearest hundreth
Answer by stanbon(75887)

what is 2a^2-30a+108=0 rounded to the nearest hundreth
------
2[a^2-15+54] = 0
----
(a-9)(a-6) = 0
-------
a = 9 or a = 6
====================
Cheers,
Stan H.
-----------------